OpenStudy (anonymous):

Sharmi ran 3 miles in 24 minutes. How long would it take her to run 7 miles? choices: 56 minutes 28 minutes 48 minutes 21 minutes

Parth (parthkohli):

= 1 mile in 8 minutes(24/3 = 8) So, multiply 8 and 7 and you get 56 minutes as the answer

OpenStudy (anonymous):

can you help me with this last question please!! What value of x makes the ratios equivalent? x/6 = 18/12 choices: 9 2 4 6

Parth (parthkohli):

Okay.. 18 * 6 = 12x => 108 = 12x => x = 108/12 => x = 9

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!!!!!!!!!!!